OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. 8 + 27 + 64 + 125 + ... + n3

OpenStudy (anonymous):

so tell me - what is the first n ?

OpenStudy (anonymous):

we know that the first n will be 2 since 2^3 = 8 so the summation start with n =2 and goes to infinity and we sum for n^3 : \[\sum_{n=2}^{\infty} n^3\]

OpenStudy (anonymous):

so as you can see at the bottom we write the index n and its initial value 2 at the top we write the final value for n which is infinity in our case since it wont end (continues) inside the summation notation we write the expression - in our case n^3

OpenStudy (anonymous):

so you just plug the number into the n^3?

OpenStudy (anonymous):



OpenStudy (anonymous):

i hope it answers the question

OpenStudy (anonymous):

Yes i understand thank you! How would you do one that doesn't have an "n" like -1 + 2 + 5 + 8 + ... + 44 <<< this one

OpenStudy (anonymous):

so we need to come up with a formula that will make our sequence you see that the difference between two elements is 3 so we can write something like : -1 + 3n and start from n=0 you can see now that when n=0 we get -1 and when n=1 we get 2 etc.. we will sum to the n that will give us 44 (what n is it ?)

OpenStudy (anonymous):

15??

OpenStudy (anonymous):

and did you get the -1 in the equation just because its the first one in the sequence?

OpenStudy (anonymous):

15 - correct

OpenStudy (anonymous):

and for the -1 : we could write the formula -4 + 3n as well and start with n=1 instead of n=0 (and end with n =16 instead of our 15)

OpenStudy (anonymous):

we could as well write 2+3n and start with n=-1 and end with n =14

OpenStudy (anonymous):

but it is easier to start with n=0 and have the first term in the sequence in your formula

OpenStudy (anonymous):

ok thank you soooo much! (:

OpenStudy (anonymous):

yw :)